OpenStudy (anonymous):

Solve the exponential equation using "relating the bases" Sq rt 5=125^x

myininaya (myininaya):

\[5^\frac{1}{2}=125^x\] \[5^\frac{1}{2}=5^{3x}\] so 1/2=3x so 1/2*1/3=x x=1/6